OpenStudy (anonymous):

the number of calories you burn depend on your weight. A 140- pound person burns 207 calories during 30 mins of tennis. Find the number of calories that a 240 pound person would burn during the same activity. How many calories would a 240- pound person burn?

OpenStudy (anonymous):

If a 140 lb person burns 207 calories, then a 240 lb person burns x calories proportion time.... 140/207 = 240/x (cross multiply) 140x = (207)(240) 140x = 49680 x = 49680/140 x = 354.8 when rounded equals 355.